# Structured Learning of Compositional Sequential Interventions

### v1

@@ Abstract
We consider sequential treatment regimes where each unit is exposed to combinations of interventions over time. When interventions are described by qualitative labels, such as “close schools for a month due to a pandemic” or “promote this podcast to this user during this week”, it is unclear which appropriate structural assumptions allow us to generalize behavioral predictions to previously unseen combinatorial sequences. Standard black-box approaches mapping sequences of categorical variables to outputs are applicable, but they rely on poorly understood assumptions on how reliable generalization can be obtained, and may underperform under sparse sequences, temporal variability, and large action spaces. To approach that, we pose an explicit model for composition, that is, how the effect of sequential interventions can be isolated into modules, clarifying which data conditions allow for the identification of their combined effect at different units and time steps. We show the identification properties of our compositional model, inspired by advances in causal matrix factorization methods but focusing on predictive models for novel compositions of interventions instead of matrix completion tasks and causal effect estimation. We compare our approach to flexible but generic black-box models to illustrate how structure aids prediction in sparse data conditions.
@@ 1 Contribution
Figure 1: Within unit $n$, interventions $D_{n}^{t}$ interact with (latent) random effect parameters $\beta_{n}$ to produce behaviour $X_{n}^{t}$ represented as a dense graphical model with square vertices denoting interventions (Dawid, 2021). Further assumptions will be required for the identifiability of the impact of interventions and their combination, including temporal impact.
Many causal inference questions involve a treatment sequence that varies over time. In the (discrete) time-varying scenario, an unit $n$ is exposed to a sequence $D_{n}^{1},D_{n}^{2},D_{n}^{3},\dots$ of treatments controlled by an external agent (Hernán and J, 2020, Ch. 19). Each $D_{n}^{t}$ can be interpreted as a cause of the subsequent behavior $X_{n}^{t},X_{n}^{t+1},\dots$, and its modeling amenable to the tools of causal inference (e.g., Robins, 1986, 1997; Murphy, 2003; Didelez et al., 2006; Chakraborty and Moodie, 2013; Zhang and Bareinboim, 2020). For instance, each $D_{n}^{t}$ can take values in the space of particular drug dosages which can be given to an in-patient, or combinations of items to be promoted to a user of a recommender system. Here, patients and users are the units, and $X_{n}^{t}$ denotes a measure of their state or behavior under time $t$. Dependencies can result in a dense causal graph, as in Fig. 1. Further background on this classical problem is provided in Appendix A.
@@
We introduce an approach to model the effects of a combinatorial and sparse space of intervention sequences, including the bursts of non-stationarity that the application of an intervention brings in. We consider the case where each $D_{n}^{t}$ behaves as if it was randomized to focus on predictive guarantees on large, possibly underspecified, sequential action spaces. We will consider neither experimental design nor reinforcement learning, instead focusing on the problem of behavioral forecasting under hypothetical future interventions given past controlled interventions.
Challenges and prior work.
@@
Much of the sequential intervention literature in AI and statistics considers generic models [^1] We include here models which explicit parameterize causal contrasts in longitudinal interventions, such as the blip models of Robins (1997). to map sequences $D_{n}^{1:t}:=[D_{n}^{1},D_{n}^{2},\dots,D_{n}^{t}]$ to behavior $X_{n}^{t}$ and beyond, sometimes exploiting strong Markovian assumptions or short sequences and small action spaces (Chakraborty and Moodie, 2013; Sutton and Barto, 2018). An unstructured black-box for sequences, such as recurrent neural networks and their variants (e.g. Hochreiter and Schmidhuber, 1997; Cho et al., 2014; Vaswani et al., 2017), can learn to map a sequence to a prediction. This is less suitable when interventions are categorical and sparse, with most entries corresponding to some baseline category of “no changes”. That is, there is no obvious choice of smoothing or interpolation criteria to generalize from seen to unseen strings $D_{n}^{1:T}$, with the practical assumption being that we see enough variability to reliably perform this interpolation. There is an increased awareness of problems posed by large discrete spaces among pre-treatment covariates (Yao et al., 2019; Zeng et al., 2024), which is of a different nature as losses can be minimized with respect to an existing natural distribution, as opposed to the interventional problem where the design variables are allowed to be chosen and so sparsity conditions can easily result in very atypical test cases of interest. Several papers address large intervention spaces (mostly for non-sequential problems) without particular concerns about non-smooth, combinatorial identifiability (Kaddour et al., 2021; Nabi et al., 2022; Saito et al., 2023, 2024). A few more recent papers address explicit assumptions of identifiability directly in combinatorial categorical spaces by sparse regression (Agarwal et al., 2023a) or energy functions Bravo-Hermsdorff et al. (2023), but without a longitudinal component. More classically, tools like the do-calculus (Pearl, 2009) and their extensions (Correa and Bareinboim, 2020) can be used with directed acyclic graphs (DAGs) to infer combinatorial effects, although they often restrict each intervention variable to target a single random variable (e.g. Aglietti et al. (2020)). With strong assumptions and computation, this may include latent variables Zhang et al. (2023). Alternative representation learning ideas for carefully extrapolating to unseen interventions are presented by (Saengkyongam et al., 2024) in the non-sequential setting.
@@
In this paper, we consider the conservative problem of identification guarantees for the effect of sequential combinations of interventions which may never have co-occurred jointly in training data, formalizing assumptions that allows for the transfer and recombination of information learned across units. These are not of the same nature as identifiability of causal effects from observational data (Pearl, 2009) or causal discovery (Spirtes et al., 2000), but of causal extrapolation: even in the randomized (or unconfounded) regime, it is not a given that the distribution of a particular potential outcome $X(d^{1:t})$ can be reliably inferred from limited past combinations of interventions without an explicit structure on the causal generative model. We will consider the case where interventions have non-stationary effects: once unit $n$ is exposed to $D_{n}^{t}=d_{n}^{t}$, then $X_{n}^{t+\Delta}$ will have dynamics that can be affected not only by the choice of $d_{n}^{t}$ but also the difference $\Delta$, resulting in dense dependency structures such as the one in Figure 1. As multiple treatments are applied to the same unit $n$, there will be a composition of effects that will depend not only on the labels of the interventions, but also on the order by which they take place. If a substantive number of choices of interventions at time $t$ is available, a large dataset may have no two sequences $D_{n}^{1:T}$ and $D_{n^{\prime}}^{1:T}$ taking the same values, even when considering only the unique treatment values applied $n$ and $n^{\prime}$ regardless of ordering and time stamps.
@@
Let each individual unit $n$ be described by time-invariant features $Z_{n}$ and a pair of time-series ($(\{X_{n}^{t}\},\{D_{n}^{t}\})$. Each $D_{n}^{t}$ is a categorical variable representing an intervention (also called treatment, or exposure), encoded by values in $\mathbb{N}$, and sequentially unconfounded with the system by randomization or assumption. $D_{n}^{t}$ describes an intervention which causes changes to the distribution of $X_{n}^{t:\infty}$, the time-series of state variables starting at time $t$. The special value $D_{n}^{t}=0$ denotes an “idle” or “default” treatment that can be interpreted as “no intervention”. Given a training set of units $(Z_{n},\{X_{n}^{t}\},\{D_{n}^{t}\})$ and observations up to a time point $T$, the goal is to predict future sequence of states for an unit $n^{\star}$ under a hypothetical intervention $D_{n^{\star}}$. We denote it as $X_{n^{\star}}^{T+1:T+\Delta}(d_{n^{\star}}^{1:T+\Delta})$, where $V^{t_{1}:t_{2}}$ denotes a slice of a discrete-time time series of random vectors $V$ from $t_{1}$ to $t_{2}$, inclusive; $V(d^{1:t})$ is the potential outcome (Imbens and Rubin, 2015; Hernán and J, 2020; Pearl, 2009) of variable $V$ under interventions $D^{1}=d^{1},D^{2}=d^{2},...,D^{t}=d^{t}$. This process takes place under stable unit treatment value assumption (SUTVA) (Imbens and Rubin, 2015), meaning that interventions $D_{n}^{t}$ only affect unit $n$.
Each intervention $D_{n}^{t}$ is allowed to have an impact on the whole future series $t,t+1,t+2,\dots$. It can correspond to an instantaneous shock (“give five dollars of credit at time $t$”) or an action that takes place over an extended period of time (“promote this podcast from time $t$ to time $t+5$”): we just assume the meaning to be directly encoded within arbitrary category labels $0,1,2,3,\dots$. In a real-world scenario where a same action can be applied more than once, different symbols in $\mathbb{N}$ should be used for each instance (e.g., 1 arbitrarily standing for “give five dollars”, and 2 for “give five dollars a second time” and so on).
@@ 2 A Structural Approach for Intervention Composition
For simplicity, we will assume scalar behavioral measurements $X_{n}^{t}$, as the multivariate case readily follows. Consider the following conditional mean model for the potential outcome $X_{n}^{t}(d_{n}^{1:t})$,
$\mathbb{E}[X_{n}^{t}(d_{n}^{1:t})~{}|~{}x_{n}^{1:t-1}(d_{n}^{1:t-1}),z_{n}]=(\phi_{n}^{t})^{\mathsf{T}}(\beta_{n}\odot\psi_{n}^{t})=\sum_{l=1}^{r}\phi_{nl}^{t}\beta_{nl}\psi_{nl}^{t},$ (1)
where $\odot$ is the elementwise product, $X_{n}^{t}(d_{n}^{1:t^{\prime}})=X_{n}^{t}(d_{n}^{1:t})$ for any $t^{\prime}\geq t$ (future interventions do not affect past outcomes), and $Z_{n}(d_{n}^{1:t})=Z_{n}$ for all $t$. We do not explicitly condition on $D_{n}^{1:t}$ in most of our notation, adopting a convention where potential outcome indices $d_{n}^{1:t}$ are always in agreement with the (implicit) corresponding observed $D_{n}^{1:t}$. We define $\phi_{nl}^{t}:=\phi_{l}(x_{n}^{1:t-1}(d_{n}^{1:t-1}),z_{n})$ as evaluations of basis functions $\phi_{l}(\cdot)$, which for now we will assume to be known. Each individual $n$ has their individual-level coefficients $\beta_{n}$. Causal impact, as attributed to $D_{n}^{1:t}=d_{n}^{1:t}$, is given by
@@
where function $\psi_{l}(d,t^{\prime},t)$ captures a trajectory in time for the effect of $D_{n}^{t^{\prime}}=d$, for $t^{\prime}<t$. We consider two variants. The first is a time-bounded variant which, for $d>0$, defines $\psi_{l}(d,t^{\prime},t):=w_{dl,t-t^{\prime}}$, where $w_{dl,t-t^{\prime}}$ is a free parameter for $0\leq t-t^{\prime}<k_{d}$, otherwise $w_{dl,t-t^{\prime}}=w_{dl,k_{d}-1}$. The intuition is that the effect of intervention level $d$ has no shape constraints but it cannot affect the past and it must settle to a constant after a chosen time window hyperparameter $k_{d}>0$. The default level $d=0$ has no effect and no free parameters, with $\psi_{l}(0,t^{\prime},t)=1$. The second variant is motivated by real-world phenomena where causal impacts diminish their influence over time and result in a new equilibrium (Brodersen et al., 2015). We consider a time-unbounded shape with exponential decay, parameterized as
@@
where $\sigma(\cdot)$ is the sigmoid function, so that $10<\sigma(w_{1d})<$ 1. As $t$ grows, $\sigma(w_{1d})^{t-t^{\prime}}$ goes to 0. $w_{3d}$ is the stationary contribution of an intervention at level $d$. In particular, for $d=0$, we define $w_{10l}=w_{20l}=0$ and $w_{30l}=1$, while for $d>0$ we have that $w_{1dl}$, $w_{2dl}$ and $w_{3dl}$ are free parameters to be learned. By defining $k_{d}$ as the dimensionality of intervention level $d$, we adopt the convention $k_{d}=3$ for the prescribed time-unbounded model of Eq. (3).
2.1 Representation Power and Compositional Warping

### v2

@@ Abstract
We consider sequential treatment regimes where each unit is exposed to combinations of interventions over time. When interventions are described by qualitative labels, such as “close schools for a month due to a pandemic” or “promote this podcast to this user during this week”, it is unclear which appropriate structural assumptions allow us to generalize behavioral predictions to previously unseen combinations of interventions. Standard black-box approaches mapping sequences of categorical variables to outputs are applicable, but they rely on poorly understood assumptions on how reliable generalization can be obtained, and may underperform under sparse sequences, temporal variability, and large action spaces. To approach that, we pose an explicit model for composition, that is, how the effect of sequential interventions can be isolated into modules, clarifying which data conditions allow for the identification of their combined effect at different units and time steps. We show the identification properties of our compositional model, inspired by advances in causal matrix factorization methods. Our focus is on predictive models for novel compositions of interventions instead of matrix completion tasks and causal effect estimation. We compare our approach to flexible but generic black-box models to illustrate how structure aids prediction in sparse data conditions.
@@ 1 Contribution
Figure 1: Within unit $n$, actions $D_{n}^{t}$ interact with (latent) random effect parameters $\beta_{n}$ to produce behavior $X_{n}^{t}$ represented as a dense graphical model with square vertices denoting interventions $\mathrm{do}(d_{n}^{1:t})$ (Pearl, 2009; Dawid, 2021). Further assumptions will be required for the identifiability of the impact of interventions and their combination, including how temporal impact takes shape and the number of independent units of observation.
Many causal inference questions involve a treatment sequence that varies over time. In the (discrete) time-varying scenario, an unit $n$ is exposed to a sequence $D_{n}^{1},D_{n}^{2},D_{n}^{3},\dots$ of actions, or treatments, potentially controllable by an external agent (Hernán and J, 2020, Ch. 19) intervening on them. Each $D_{n}^{t}$ can be interpreted as a cause of the subsequent behavior $X_{n}^{t},X_{n}^{t+1},\dots$, and its modeling amenable to the tools of causal inference (e.g., Robins, 1986, 1997; Murphy, 2003; Didelez et al., 2006; Chakraborty and Moodie, 2013; Zhang and Bareinboim, 2020). For instance, each $D_{n}^{t}$ can take values in the space of particular drug dosages which can be given to an in-patient, or combinations of items to be promoted to a user of a recommender system. Here, patients and users are the units, and $X_{n}^{t}$ denotes a measure of their state or behavior under time $t$. Dependencies can result in a dense causal graph, as in Fig. 1. Further background on this classical problem is provided in Appendix A.
@@
We introduce an approach to predict the effects of actions within a combinatorial and sparse space of categorical intervention sequences. That is, for an unit $n$ and given past behavioral sequence $X_{n}^{1:t-1}:=[X_{n}^{1},\dots,X_{n}^{t-1}]$, manipulable variables $D_{n}^{1:t-1}:=[D_{n}^{1},\dots,D_{n}^{t-1}]$ and time-fixed covariates $Z_{n}$, we predict future behavior $X_{n}^{t},X_{n}^{t+1},\dots,X_{n}^{t+\Delta}$ under the (partially hypothetical) interventional regime $\mathrm{do}(d_{n}^{1:t+\Delta})$. The latter denotes controlling the action variables by external intervention Pearl (2009). For mean-squared error losses, prediction means learning $\mathbb{E}[X_{n}^{t+\delta}~{}|~{}x_{n}^{1:t-1},z_{n},\mathrm{do}(d_{n}^{1:t+\delta})]$, $0\leq\delta\leq\Delta$. We consider the case where sequential ignorability holds, by randomization or adjustment (Pearl, 2009; Hernán and J, 2020), in order to focus on predictive guarantees for previously unseen action sequences.
Challenge.
@@
Much of the sequential intervention literature in AI and statistics considers generic black-box or parametric models [^1] We include here models which explicitly parameterize causal contrasts in longitudinal interventions, such as the blip models of Robins (1997). to map sequences $D_{n}^{1:t}:=[D_{n}^{1},D_{n}^{2},\dots,D_{n}^{t}]$ to behavior $X_{n}^{t}$ and beyond, sometimes exploiting strong Markovian assumptions or short sequences and small action spaces (Chakraborty and Moodie, 2013; Sutton and Barto, 2018). An unstructured black-box for sequences, such as recurrent neural networks and their variants (e.g. Hochreiter and Schmidhuber, 1997; Cho et al., 2014; Vaswani et al., 2017), can learn to map a sequence to a prediction. This is less suitable when actions are categorical and sparse, in the sense that most entries correspond to some baseline category of “no changes”. That is, there is no obvious choice of smoothing or interpolation criteria to generalize from seen to unseen strings $D_{n}^{1:T}$, with the practical assumption being that we observe enough variability to reliably perform this interpolation. There is an increased awareness of problems posed by large discrete spaces among pre-treatment covariates (Yao et al., 2019; Zeng et al., 2024), which is of a different nature as losses can be minimized with respect to an existing natural distribution, as opposed to the interventional problem where the action variables are allowed to be chosen and so sparsity conditions can easily result in very atypical test cases of interest. Several papers address large intervention spaces (mostly for non-sequential problems) without particular concerns about non-smooth, combinatorial identifiability (Kaddour et al., 2021; Nabi et al., 2022; Saito et al., 2023, 2024). A few more recent papers address explicit assumptions of identifiability directly in combinatorial categorical spaces by sparse regression (Agarwal et al., 2023a) or energy functions Bravo-Hermsdorff et al. (2023), but without a longitudinal component. More classically, tools like the do-calculus (Pearl, 2009) and their extensions (Correa and Bareinboim, 2020) can be used with directed acyclic graphs (DAGs) to infer combinatorial effects, although they often restrict each manipulated variable to directly affect only a small set of variables (e.g. Aglietti et al. (2020)). With strong assumptions and computational cost, this may include latent variables Zhang et al. (2023). Alternative representation learning ideas for carefully extrapolating to unseen interventions are presented by (Saengkyongam et al., 2024) in the non-sequential, non-combinatorial, setting.
@@
In this paper, we consider the conservative problem of identification guarantees for the effect of sequential combinations of interventions which may never have co-occurred jointly in training data, formalizing assumptions that allow for the transfer and recombination of information learned across units. These are not of the same nature as identifiability of causal effects from observational data (Pearl, 2009) or causal discovery (Spirtes et al., 2000), but of causal extrapolation: even in the randomized (or unconfounded) regime, it is not a given that the distribution under a particular regime $\mathrm{do}(d_{n}^{1:t})$ can be reliably inferred from limited past combinations of action sequences without an explicit structure on the causal generative model. We will consider the case where actions imply non-stationary transitions: once unit $n$ is exposed to $D_{n}^{t}=d_{n}^{t}$, then $X_{n}^{t+\Delta}$ will have transition dynamics that can be affected not only by the choice of $d_{n}^{t}$ but also the difference $\Delta$, resulting in dense dependency structures such as the one in Figure 1. As multiple treatments are applied to the same unit $n$, there will be a composition of effects that will depend not only on the labels of the actions, but also on the order by which they take place. If a substantive number of choices of treatments at time $t$ is available, a large dataset may have no two sequences $D_{n}^{1:T}$ and $D_{n^{\prime}}^{1:T}$ taking the same values, even when considering only the unique treatment values applied to $n$ and $n^{\prime}$, regardless of ordering and time stamps.
@@
Let each individual unit $n$ be described by time-invariant features $Z_{n}$ and a pair of time-series $(X_{n}^{1:T},D_{n}^{1:T})$. Here, $V^{t_{1}:t_{2}}$ denotes a slice of a discrete-time time-series of variables $V^{t}$ from $t_{1}$ to $t_{2}$, inclusive. That observations for all units stop at time $T$ is not a requirement, but just a way of simplifying presentation. Each $D_{n}^{t}$ is a categorical action variable which, when under external control, represents a potential intervention, encoded by values in $\mathbb{N}$. We assume data is available under a sequentially ignorable selection of actions (Pearl, 2009; Hernán and J, 2020), either by randomization or consequences of an adjustment. $D_{n}^{t}$ describes a (possibly hypothetical) intervention which causes changes to the distribution of $X_{n}^{t:\infty}$. The special value $D_{n}^{t}=0$ denotes an “idle” or “default” treatment level that can be interpreted as “no intervention”. We are given a training set of units $(Z_{n},X_{n}^{1:T},D_{n}^{1:T})$. The goal is to predict future sequence of behavior $X_{n^{\star}}$ for a (training or test) unit $n^{\star}$ under a hypothetical intervention sequence $\mathrm{do}(d_{n^{\star}}^{1:T+\Delta})$. This process takes place under the stable unit treatment value assumption (SUTVA) (Imbens and Rubin, 2015), meaning that action variables $D_{n}^{t}$ only affect unit $n$.
Each intervention on $D_{n}^{t}$ is allowed to have an impact on the whole future series $t,t+1,t+2,\dots$. It can correspond to an instantaneous shock (“give five dollars of credit at time $t$”) or an action that takes place over an extended period of time (“promote this podcast from time $t$ to time $t+5$”): we just assume the meaning to be directly encoded within arbitrary category labels $0,1,2,3,\dots$. In a real-world scenario where a same action can be applied more than once, different symbols in $\mathbb{N}$ should be used for each instance (e.g., 1 arbitrarily standing for “give five dollars”, and 2 for “give five dollars a second time” and so on).
@@ 2 A Structural Approach for Intervention Composition
For simplicity, we will assume scalar behavioral measurements $X_{n}^{t}$, as the multivariate case readily follows. Consider the following conditional mean model for the regime $\mathrm{do}(d_{n}^{1:t})$,
$\mathbb{E}[X_{n}^{t}~{}|~{}x_{n}^{1:t-1},z_{n},\mathrm{do}(d_{n}^{1:t})]=(\phi_{n}^{t})^{\mathsf{T}}(\beta_{n}\odot\psi_{n}^{t})=\sum_{l=1}^{r}\phi_{nl}^{t}\beta_{nl}\psi_{nl}^{t},$ (1)
where $\odot$ is the elementwise product, with $\phi_{n}^{t}$, $\beta_{n}$ and $\psi_{n}^{t}$ defined below. We postpone explaining the motivation for this structure to Section 2.1. First, we describe this structure in more detail.
Using $p(\cdot)$ to denote generic probability mass/density functions for random variables given by context, assume $p(x_{n}^{t}~{}|~{}\mathrm{do}(d_{n}^{1:t^{\prime}}))=p(x_{n}^{t}~{}|~{}\mathrm{do}(d_{n}^{1:t}))$ for any $t^{\prime}\geq t$ (future interventions do not affect the past), and $p(z_{n}~{}|~{}\mathrm{do}(d_{n}^{1:t}))=p(z_{n})$ for all $t$. We do not explicitly condition on $D_{n}^{1:t}$ in most of our notation, adopting a convention where intervention indices $\mathrm{do}(d_{n}^{1:t^{\prime}})$ are always in agreement with the (implicit) corresponding observed $D_{n}^{1:t}=d_{n}^{1:t}$. We define $\phi_{nl}^{t}:=\phi_{l}(x_{n}^{1:t-1},z_{n})$ as evaluations of basis functions $\phi_{l}(\cdot)$. For now, we will assume such $\phi_{l}(\cdot)$ functions to be known. Each unit $n$ has individual-level coefficients $\beta_{n}$. Causal impact, as attributed to $\mathrm{do}(d_{n}^{1:t})$, is given by
@@
where function $\psi_{l}(d,t^{\prime},t)$ captures a trajectory in time for the effect of $D_{n}^{t^{\prime}}=d$, for $t^{\prime}<t$. Compositionality will be described in the sequel in two ways: first, as a way of combining the effects of actions in a type of recursive functional composition (compositional recursion); second, by interpreting this function composition at time $t$ in terms of an operator warping a baseline unit-level feature vector $\beta_{n}$ with an interventional-level embedding vector $\psi(d_{n}^{t},t,t)$ (compositional warping).
Effect families.
We consider two variants for $\psi(\cdot)$. The first is a time-bounded variant which, for $d>0$, defines $\psi_{l}(d,t^{\prime},t):=w_{dl,t-t^{\prime}}$, where $w_{dl,t-t^{\prime}}$ is a free parameter for $0\leq t-t^{\prime}<k_{d}$, otherwise $w_{dl,t-t^{\prime}}=w_{dl,k_{d}-1}$. The intuition is that the effect of intervention level $d$ has no shape constraints but it cannot affect the past and it must settle to a constant after a chosen time window hyperparameter $k_{d}>0$. The default level $d=0$ has no effect and no free parameters, with $\psi_{l}(0,t^{\prime},t)=1$. The second variant is motivated by real-world phenomena where causal impacts diminish their influence over time and result in a new equilibrium (Brodersen et al., 2015). We consider a time-unbounded shape with exponential decay, parameterized as
@@
where $\sigma(\cdot)$ is the sigmoid function, so that $10<\sigma(w_{1d})<$ 1. As $t$ grows, $\sigma(w_{1d})^{t-t^{\prime}}$ goes to 0. $w_{3d}$ is the stationary contribution of an intervention at level $d$. In particular, for $d=0$, we define $w_{10l}=w_{20l}=0$ and $w_{30l}=1$, while for $d>0$ we have that $w_{1dl}$, $w_{2dl}$ and $w_{3dl}$ are free parameters to be learned. Let us interpret the previous $k_{d}$ hyperparameter as the dimensionality of intervention level $d$. As the model of Eq. (3) is given by vectors $w_{1d},w_{2d},w_{3d}$, we here define $k_{d}=3$.
2.1 Representation Power: Compositional Recursion and and Compositional Warping